OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). (2 points) Question 1 options: 1) y = x + 7 2) y = -3x + 3 3) y = x + 3 4) y = -3x + 7

OpenStudy (texaschic101):

y = -3x + 12 first we need to find the slope in the equation given. In y = mx + b form, the m position is the slope. Therefore, the slope(m) = -3. A parallel line will have the same slope. So the slope we need will also be -3. Now we use y = mx + b, and since we already know the slope(m), we need to find the y intercept(b). y = mx + b slope(m) = -3 points given : (-1,6)....x = -1 and y = 6 now we sub and solve for b, the y intercept 6 = -3(-1) + b 6 = 3 + b 6 - 3 = b 3 = b so our parallel line will be : y = -3x + 3

OpenStudy (anonymous):

thank you!! do you have time to answer one more for me please?